OpenStudy (anonymous):

Did I do this Geometry problem correctly? I think it's 7? (je crois que c'est sept)

OpenStudy (anonymous):

http://img845.imageshack.us/img845/6218/danish1.jpg

OpenStudy (anonymous):

did you compute \(\sqrt{13^2-5^2}\) ?

OpenStudy (anonymous):

Oh, it's 12! Silly me! :P

OpenStudy (anonymous):

Care to do another another one? :)

OpenStudy (anonymous):

\(x\) cannot be 7 because that would mean it the sum of the legs of the triangle was shorter than the hypotenuse!

OpenStudy (anonymous):

sure if i can my geometry is not that good

OpenStudy (anonymous):

That's fine! :)

OpenStudy (anonymous):

http://img843.imageshack.us/img843/9726/danish2.jpg

OpenStudy (anonymous):

I think it's 34 inches

OpenStudy (kinggeorge):

I'm getting 34 as well.

OpenStudy (anonymous):

Oh thank you! Let's continue.

OpenStudy (anonymous):

yeah i like 34 whew, my geometry is very weak

OpenStudy (anonymous):

http://img52.imageshack.us/img52/2969/danish3.jpg

OpenStudy (anonymous):

I think it's 108

OpenStudy (anonymous):

If you're uncertain, do you think you could send a Geometry expert over here? :)

OpenStudy (anonymous):

AccessDenied! You're reading this! :D

OpenStudy (anonymous):

Aaand he's gone

OpenStudy (anonymous):

Oh, thank you! Could you stay for some more, eliass?

OpenStudy (anonymous):

Ask it as a new problem s'il te plait

OpenStudy (anonymous):

Wait a minute...it's not 120...

OpenStudy (anonymous):

re-do the problem, please

OpenStudy (kinggeorge):

I'm getting 105. You have 45 from BCA, and since arc CE is 50, you know arc EA is 180-50=130. Thus, angle ACE is 130/2=65. Thus, angle BCE=45+65=105.

OpenStudy (anonymous):

65+45=110

OpenStudy (kinggeorge):

Derp. It is 110.

OpenStudy (anonymous):

ok ha ha KingGeorge, can you stay for a bit? :)

OpenStudy (kinggeorge):

I'm going to agree with elias. Please ask it as a new question.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

http://img208.imageshack.us/img208/5784/danish4.jpg

OpenStudy (anonymous):

it's 12 right?